OpenStudy (anonymous):

how would y=3x+11 be turned into standard form?

OpenStudy (anonymous):

3x+y=11

OpenStudy (anonymous):

and the answer is 3x+7y=11

OpenStudy (anonymous):

can you explain to me how you got that answer?

Directrix (directrix):

The Standard Form for a linear equation in two variables, x and y, is usually given as Ax + By = C where, if at all possible, A, B, and C are integers, and A is non-negative, and, A, B, and C have no common factors other than 1.

OpenStudy (anonymous):

@Directrix i understand, however i dont get how to convert the two equations

OpenStudy (anonymous):

nvm, i submitted the assignment. thank you.

Directrix (directrix):

>>how would y=3x+11 be turned into standard form? y=3x+11 y = 3x + 11 3x + 11 = y 3x - y + 11 = y - y 3x - y + 11 = 0 Now, @jocelynsalcedo subtract 11 from both sides.

Directrix (directrix):

I think Baby Slim may have a typo here with the seven: > and the answer is 3x+7y=11

OpenStudy (anonymous):

well thankyou.

Directrix (directrix):

Do you want to finish the problem to determine the correct answer?